I need help on solving -5w+8>=13 and the steps to solving it.
Answer by philline_palana(20)

-5w+8>=13
-5w>=13-8
-5w>=5
-5w/-5>=5/-5
*we should change the sign into "<=" because we divide the equation by a negative number so our answer is:
w<=-1